OpenStudy (anonymous):

Calc Question. Will Medal! Evaluate the problem below. I know for a fact the answer is ex + c. But what are the steps?

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ e \sqrt{x} }{ sqrt{x} }\]

OpenStudy (anonymous):

Sorry for the format

zepdrix (zepdrix):

\[\large\rm \int\limits \frac{e^{\sqrt x}}{\sqrt x}dx\quad=\quad \int\limits e^{\sqrt x}\left(\frac{1}{\sqrt x}dx\right)\]

zepdrix (zepdrix):

Making the substitutiion, \(\large\rm u=\sqrt x\)

OpenStudy (anonymous):

Ah ok. So then du = \[\frac{ 2 }{ 3 }x ^{\frac{ 3 }{ 2 }}\]

zepdrix (zepdrix):

Oh you integrated by mistake.

zepdrix (zepdrix):

You want to `differentiate` to find your du.

OpenStudy (anonymous):

*facepalm* Sorry!

OpenStudy (anonymous):

du = \[\frac{ 1 }{ 2 }x ^{\frac{ -1 }{ 2 }}\]

zepdrix (zepdrix):

Bahhh you can do your 1/2 power if you want. But square root comes up so often that you should just put this derivative to memory,\[\large\rm \frac{d}{dx}\sqrt{x}\quad=\quad \frac{1}{2\sqrt x}\]Derivative of square root is one over two square roots.

OpenStudy (anonymous):

Good to know! Will do!

zepdrix (zepdrix):

\[\large\rm du=\frac{1}{2\sqrt x}dx\]Hmm this is ALMOST exactly what we bracketed in our integral. Is there any way we can get rid of the 2?

OpenStudy (anonymous):

Put it outside the integral, making it look like: \[2\int\limits_{?}^{?}\frac{ 1 }{ \sqrt{x} }\]

zepdrix (zepdrix):

Well, yes but I meant in our substitution, not in the integral. You can't just make a 2 magically appear like that :)\[\large\rm du=\frac{1}{2\sqrt x}dx\]Multiplying both sides by 2 gives us,\[\large\rm 2du=\left(\frac{1}{2\sqrt x}dx\right)\]

zepdrix (zepdrix):

\[\large\rm \int\limits\limits e^{\sqrt x}\left(\frac{1}{\sqrt x}dx\right)\quad=\quad \int\limits e^u\left(2du\right)\]

zepdrix (zepdrix):

Ahh I made a typo :(

OpenStudy (anonymous):

I see what your saying!

zepdrix (zepdrix):

Multiplying by 2 gave us,\[\large\rm 2du=\left(\frac{1}{\sqrt x}dx\right)\]

OpenStudy (anonymous):

Ok! So knowing this, we can now make it so that the 2 is upfront and the function left (e^sqrt(x)). So now would we just have to integrate?

zepdrix (zepdrix):

\[\large\rm =2\int\limits e^u du\]Yes, we can integrate from here.

OpenStudy (anonymous):

Awesome! So the final answer would be \[2e ^{u}\] and substituting in the property of u, making it \[2e ^{\sqrt{x}} + C\]

zepdrix (zepdrix):

Yes, good job!

OpenStudy (anonymous):

That was similar than I thought! Haha thank you!

OpenStudy (anonymous):

*simpler

zepdrix (zepdrix):

heh :) np